Plan: to measure Fe lantanides, simultaneous Ka and Kb acquisition; users are interested in HERFD on Kbeta. Not possible to measure V2C simultaneously with Ka; can be later after shifting the detector.
Ka: Si(333) 67.8 o.
Kb: Si(440) 66.2 o.
Further users are potentially interested to mount the cryo and measure Se and Co spectra.
Took until ~ 6 pm to align. Problems with ZMX crate (it broke on the weekend).
Users measured Fe references from ~ 8 pm.
Kalpha, Kbeta and HERFD during the day and for the night.
High harmonics were causing some extra reflections; were removed by going away from the middle of the pitch to ~90% of intensity.
The cryo was installed during the day.
Ka/Kb in the same spectrometer configuration was measured in the evening and night. Alignment was not easy, because samples were not thin: there was sometimes a double layer.
Next set of samples go into the cryo at 2 pm. This time they are supposed to be finer powder and the captons should be stuck together.